Phantomdex:

An exponential model is given to be ŷ = 3.5(2)x. Determine the predicted value when x = 1.6. 22.499 10.610 7.0 5.6

Lui0210:

Okay so for this one, you can just plug it into a calculator as 3.5(2)^(1.6)

Lui0210:

Which should give you ~10.61002

Lui0210:

Thunderstruck:

C. 7.0

RhianaB112:

11.2 Sure! Let's substitute the given value of xx into the equation \hat{y} = 3.5(2)^x y ^ ​ =3.5(2) x . \hat{y} = 3.5(2)^{1.6} y ^ ​ =3.5(2) 1.6 \hat{y} = 3.5(3.2) y ^ ​ =3.5(3.2) \hat{y} = 11.2 y ^ ​ =11.2 Therefore, the predicted value when x=1.6x=1.6 is 11.2.